There are 5 houses on a street namely, A, B, C, D and E. For all cases, assume that positions to the right are positive and distance between each house is 10 m.

a) Draw a frame of reference with house 'A' as the original position.

b) If you live in house 'C', what is your position relative to house 'E'?

c) What are the positions of house 'A' and 'D', if house 'B' is taken as the reference point?
1
Expert's answer
2020-05-11T20:16:21-0400

a)




b)


"x_C-x_E=20-40=-20\ m"

c)


"x_A-x_B=0-10=-10\ m\\x_D-x_B=30-10=20\ m"